# Unlearning Reveals the Influential Training Data of Language Models

### v1

@@ Abstract
In order to enhance the performance of language models while mitigating the risks of generating harmful content, it is crucial to identify which training dataset affects the model’s outputs. Ideally, we can measure the influence of each dataset by removing it from training; however, it is prohibitively expensive to retrain a model multiple times. This paper presents UnTrac, which estimates the influence of a training dataset by unlearning it from the trained model. UnTrac is extremely simple; each training dataset is unlearned by gradient ascent, and we evaluate how much the model’s predictions change after unlearning. We empirically examine if our methods can assess the influence of pretraining datasets on generating toxic, biased, and untruthful content. Experimental results demonstrate that our method estimates their influence much more accurately than existing methods while requiring neither excessive memory space nor multiple model checkpoints.
Unlearning Reveals the Influential Training Data of Language Models
@@ 1 Introduction
Large language models (LLMs) have had a significant impact on our society. They exhibit remarkable abilities (e.g., chain-of-thought reasoning) without being explicitly trained for such tasks. At the same time, LLMs also pose potential risks, such as the amplification of discrimination through the propagation of toxic language. LLMs are trained on a vast number of corpora via pretraining or refined through finetuning on diverse tasks. Although some efforts have been made to unravel the black box of LLMs (e.g., Han et al., 2023; Wang et al., 2023a), it is still unclear which data sources cause their unprecedented abilities and potential harms.
Ideally, we can answer this question by removing each dataset from the training datasets and assessing the change in the model’s performance after retraining (leave-dataset-out). However, since we need to retrain a model on each dataset, leave-dataset-out is prohibitively expensive. Training data attribution methods overcome this problem by approximating the influence with Hessian-based influence functions (HIF; Koh and Liang, 2017; Koh et al., 2019) or tracking changes in test loss during training (TracIn; Pruthi et al., 2020). However, HIF requires a large memory footprint to approximate the inverse Hessian, while TracIn generally needs multiple model checkpoints.
@@
In our experiments, we first examine whether our methods can trace influential training tasks in the setting of finetuning. We created a dataset representing a mixture of synthetic tasks, designed to evaluate our method’s capability in assessing the influence of each task. This influence is estimated across all task pairs. In order to make this more challenging, we have created task versions that, while semantically distinct, require responses in the same format from the model. Additionally, we include pairs that are nearly identical in content but demand responses in differing formats. This approach ensures that our methods are not overly reliant on superficial similarities between tasks. We show that our methods accurately assess the influence of training tasks, where we use the expensive leave-dataset-out method as the ground-truth, and are only slightly affected by the output format.
@@ 2 Problem Formulation
$\!\mathcal{I}_{\mathrm{truth}}(\mathcal{Z}^{\prime},\mathcal{Z})$ $\!=\!\sum_{j=1}^{N^{\prime}}L(z^{\prime}_{j},\bm{\theta}_{0})\!-\!L(z^{\prime}_{j},\bm{\theta}_{-\mathcal{Z}})$ (1)
where $z^{\prime}_{j}$ is the $j$-th batch in the test dataset $\mathcal{Z}^{\prime}$, $N^{\prime}$ is the number of test batches, and $L$ is the loss function. Koh et al. (2019) use all examples in the dataset to train the counterfactual model: $\bm{\theta}_{-\mathcal{Z}}\!=\!\mathrm{arg}\min_{\bm{\theta}}\sum_{z\in\mathcal{D}\setminus\mathcal{Z}}L(z,\bm{\theta})$. This definition overly emphasizes the influence of large datasets, as, when removing them, the number of training examples drops substantially. However, when asking about the influence of a dataset, our primary interest often centers on whether the type of data present within a dataset wields considerable influence. We modify the definition of dataset’s influence so as to better align with this research question. Thus, for every training dataset $\mathcal{Z}$, we train a model for the same number of training steps $T$ instead of the entire dataset: $\bm{\theta}_{-\mathcal{Z}}=\mathrm{arg}\min_{\bm{\theta}}\sum_{t=1}^{T}L(z_{t},\bm{\theta})$ where $z_{t}\sim\mathcal{D}\setminus\mathcal{Z}$. This setup is practical for evaluating the influence of datasets of different sizes. [^1] We show the conventional setup overestimates the influence of large datasets in Appendix A.1.
@@ 3.1 UnTrac
where $T$ is the number of unlearning steps. Here, $\bm{\theta}_{i}$ is updated via gradient ascent, which maximizes the loss of the $i$-th batch $z_{i}$ in the training dataset $\mathcal{Z}$. If we use stochastic gradient ascent, the updated parameter can be written as Equation (3); however, any optimizer can be used for unlearning.
@@ 3.2 UnTrac-Inv
This alternative influence can be regarded as an approximation to UnTrac, Equation (2). Note that $\bm{\theta}_{j}=\bm{\theta}_{j-1}+\eta_{j}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{j-1})$, and use the first-order approximation $L(z_{i},\bm{\theta}_{j})-L(z_{i},\bm{\theta}_{j-1})\approx\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{j-1})(\bm{\theta}_{j}-\bm{\theta}_{j-1})$. Equation (2) and (4) can then be re-approximated as Equation (5) and (6), respectively.
@@ 4.1 TracIn, GradDot & GradCos
Pruthi et al. (2020) proposed TracIn, which traces the influence of a training example by the total change in test loss during training. While the original TracIn assesses the influence of individual training examples, it can be easily extended to assess a whole dataset. As observing the loss reduction at every step is computationally expensive, they approximate it using model checkpoints, assuming that every example in the training dataset is encountered once between these checkpoints. By approximating the loss reduction with gradients at each checkpoint $t$: $L(z^{\prime}_{j},\bm{\theta}_{t})-L(z^{\prime}_{j},\bm{\theta}_{t+1})\approx\eta_{t}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{t})^{\top}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{t})$, TracIn is defined as:
@@
where $\mathcal{T}_{cp}$ denotes the set of training steps where the model checkpoints are saved. As using multiple checkpoints induces substantial overhead, only the last checkpoint is often used in practice Schioppa et al. (2023), which is referred to as GradDot. Barshan et al. (2020) pointed out that some outlier training examples have significantly large gradients, leading to an overestimation of their influences. Therefore, normalizing the gradients (i.e., replacing the dot product with cosine similarity) can be effective, referred to as GradCos Han and Tsvetkov (2021).
@@ 4.2 Hessian-based Influence Functions
Hessian-based influence functions (HIF) are grounded in robust statistics Hampel (1974); Cook and Weisberg (1982) and were introduced to deep learning by Koh and Liang (2017). Koh et al. (2019) used HIF to assess the influence of multiple training examples, and the estimated influence correlates well with the ground-truth. Given a trained model $\bm{\theta}_{0}$, HIF estimates the influence of a training dataset $\mathcal{Z}$ on a test dataset $\mathcal{Z}^{\prime}$ as Equation (8).
@@ 4.3 Connection to UnTrac & UnTrac-Inv
GradDot, GradCos, and HIF can be viewed as special cases of UnTrac and UnTrac-Inv. Under the first-order approximation, GradDot is equivalent to UnTrac and UnTrac-Inv with a single step of unlearning using SGD, where a single batch contains all examples in the dataset. Similarly, our methods correspond to GradCos if we use RMSProp or Adam, while corresponding to HIF if we use Newton’s method as an optimizer. [^2] Detail explanations are provided in Appendix A.2. HIF can be regarded as providing an approximation to UnTrac.
@@ 5 Experiments
In this section, we evaluate how our methods can accurately measure the influence of training datasets across different model architectures (encoder-decoder and decoder-only) and training setups (pretraining and fine-tuning). [^3] The code will be publicly available after publication.
Since we would not generally have a validation set, it is impractical to tune the hyperparameters for each experiment. Thus, we set the hyperparameters based on the experimental results on Toxigen, one of the test datasets used in Section 5.2. The same hyperparameters are used across all the experiments. We use Adam Kingma and Ba (2014) with a constant learning rate of 5e-5, $\beta_{1}$ = 0.9, and $\beta_{2}$ = 0.999. As for UnTrac, we set the batch size to 1 and run unlearning for 1 epoch. As discussed in Section 3.2, UnTrac-Inv requires a small number of unlearning steps and a large batch size. Hence, we set the batch size as 256 (a single batch contains all test examples) and the number of unlearning steps (epochs) as 5. In Section 6, we discuss the hyperparameter sensitivity of our methods.
@@
HIF cannot be directly used as computing the inverse Hessian is prohibitively expensive. Following Koh and Liang (2017), we approximate the inverse Hessian by LISSA Agarwal et al. (2017), where the number of iterations is set to 10. Following Schioppa et al. (2022), we also use Arnoldi iteration with low-rank eigenvector projection. We set the number of Arnoldi iterations to $n$ = 25 and the number of eigenvectors to $\tilde{p}$ = 25 in Section 5.1 due to the memory constraints, while $n$ = 200 and $\tilde{p}$ = 100 in Section 5.2. Following the previous studies, the gradients of training datasets are normalized, and 256 training examples are used to approximate Hessian for each iteration.
Dataset
Input Text
Output Text
Test
What is the number that comes after {0}?
{1}
Train1: SimSame
Determine the number that succeeds {two}. Provide your answer in numerical form.
{3}
Train2: SimDiff
Determine the number that succeeds {one}. Provide your answer in words.
{two}
Train3: DisSame
Determine the length of ‘{problem}’. Provide your answer in numerical form.
{7}
Train4: DisDiff
Determine the length of ‘{align}’. Provide your answer in words.
{five}
Dataset
Input Text
Answer Text
Test
What letter remains in ‘{xmais}’ after extracting ‘{s}’, ‘{x}’, ‘{m}’, ‘{i}’?
{a}
Train1: SimSame
Identify the character left after removing ‘{0}’, ‘{1}’, ‘{4}’, ‘{7}’ from ‘{71b40}’.
{b}
Train2: SimDiff
Identify the character left after removing ‘{6}’, ‘{7}’, ‘{5}’, ‘{2}’ from ‘{27516}’.
{1}
Train3: DisSame
Identify the part of speech of ‘{problem}’. Select your answer with the associated letter. Choices: a. noun b. verb c. adjective d. adverb
{a}
Train4: DisDiff
Identify the part of speech of ‘{align}’. Select your answer with the associated number. Choices: 0. noun 1. verb 2. adjective 3. adverb
{1}
Table 1: Example of the synthetic dataset A (top) and B (bottom). {Strings in braces} are varied with each example.
Synthetic Dataset A Synthetic Dataset B
Train Dataset 1:SimSame 2:SimDiff 3:DisSame 4:DisDiff 1:SimSame 2:SimDiff 3:DisSame 4:DisDiff
GradDot 0.864 -0.107 0.836 -1.594 -0.561 1.730 -0.655 -0.513
GradCos 1.456 -0.508 0.291 -1.240 1.197 0.708 -1.307 -0.598
HIF (Arnoldi) 1.732 -0.594 -0.561 -0.577 1.368 0.015 0.074 -1.457
HIF (LISSA) -1.294 1.517 -0.111 -0.111 -0.600 1.708 -0.785 -0.323
TracIn -0.331 1.690 -0.443 -0.916 -0.174 1.638 -1.060 -0.404
UnTrac 1.330 0.600 -0.913 -1.018 1.492 0.175 -0.414 -1.253
UnTrac-Inv 1.688 -0.212 -0.651 -0.826 1.056 0.905 -0.716 -1.245
Ground Truth 1.416 0.462 -1.031 -0.847 1.150 0.837 -1.025 -0.962
Table 2: Influence of the training datasets estimated by each method. The average values are shown across four runs. The values are standardized for each method to normalize the range of values.
@@ 5.1 Tracing Influential Training Tasks
Dataset
We create two synthetic datasets, each containing one test dataset and four training datasets. The training datasets have the following characteristics compared with the test dataset.
  1. 1.
    Training dataset 1 (SimSame): similar task with the same output format.
  2. 2.
    Training dataset 2 (SimDiff): similar task with different output format.
  3. 3.
    Training dataset 3 (DisSame): dissimilar task with the same output format.
  4. 4.
    Training dataset 4 (DisDiff): dissimilar task with different output format.
Table 1 presents an example of the synthetic datasets A and B. Each training dataset and test dataset consists of 256 examples. The model is trained for 512 steps on the mixture of the four training datasets with a batch size of 2.

### v2

# Unlearning Traces the Influential Training Data of Language Models

@@ Abstract
Identifying the training datasets that influence a language model’s outputs is essential for minimizing the generation of harmful content and enhancing its performance. Ideally, we can measure the influence of each dataset by removing it from training; however, it is prohibitively expensive to retrain a model multiple times. This paper presents UnTrac: Unlearning Traces the influence of a training dataset on the model’s performance. UnTrac is extremely simple; each training dataset is unlearned by gradient ascent, and we evaluate how much the model’s predictions change after unlearning. Furthermore, we propose a more scalable approach, UnTrac-Inv, which unlearns a test dataset and evaluates the unlearned model on training datasets. UnTrac-Inv resembles UnTrac, while being efficient for massive training datasets. In the experiments, we examine if our methods can assess the influence of pretraining datasets on generating toxic, biased, and untruthful content. Our methods estimate their influence much more accurately than existing methods while requiring neither excessive memory space nor multiple checkpoints.
Unlearning Traces the Influential Training Data of Language Models
@@ 1 Introduction
Large language models (LLMs) have had a significant impact on our society. They exhibit remarkable abilities (e.g., chain-of-thought reasoning) without being explicitly trained for such tasks. At the same time, LLMs also pose potential risks, such as the amplification of discrimination through the propagation of toxic language. LLMs are trained on a vast number of corpora via pretraining or refined through finetuning on diverse tasks. Although some efforts have been made to unravel the black box of LLMs (e.g., Grosse et al., 2023; Feng et al., 2023), it is still unclear which data sources cause their unprecedented abilities and potential harms.
Ideally, we can answer this question by removing each dataset from the training datasets and assessing the change in the model’s performance after retraining (leave-dataset-out). However, since we need to retrain a model on each dataset, leave-dataset-out is prohibitively expensive. Training data attribution overcomes this problem by approximating the influence with Hessian-based influence functions (HIF; Koh and Liang, 2017; Koh et al., 2019) or tracking changes in test loss during training (TracIn; Pruthi et al., 2020). However, HIF requires a large memory space to approximate the inverse Hessian Schioppa et al. (2022), while TracIn generally needs multiple model checkpoints.
@@
In our experiments, we first examine whether our methods can trace influential training tasks in the setting of finetuning. We created a dataset representing a mixture of synthetic tasks, designed to evaluate our method’s capability in assessing the influence of each task. In order to make this assessment more challenging, we have created task pairs that, while semantically distinct, require responses in the same format. Additionally, we include pairs that are nearly identical in content but demand responses in differing formats. By estimating the influence across these task pairs, we verify that our methods are not overly reliant on superficial similarities between tasks. In this controlled dataset, we show that our methods accurately assess the influence of training tasks, where we use the expensive leave-dataset-out method as the ground-truth, and are only slightly affected by the output format.
@@ 2 Problem Formulation
$\!\mathcal{I}_{\mathrm{truth}}(\mathcal{Z}^{\prime},\mathcal{Z})$ $\!=\!\sum_{j=1}^{N^{\prime}}L(z^{\prime}_{j},\bm{\theta}_{-\mathcal{Z}})\!-\!L(z^{\prime}_{j},\bm{\theta}_{0})$ (1)
where $z^{\prime}_{j}$ is the $j$-th batch in the test dataset $\mathcal{Z}^{\prime}$, $N^{\prime}$ is the number of test batches, and $L$ is the loss function. Koh et al. (2019) use all examples in the dataset to train the counterfactual model: $\bm{\theta}_{-\mathcal{Z}}\!=\!\mathrm{arg}\min_{\bm{\theta}}\sum_{z\in\mathcal{D}\setminus\mathcal{Z}}L(z,\bm{\theta})$. This definition overly emphasizes the influence of large datasets, as, when removing them, the number of training examples drops substantially. However, when asking about the influence of a dataset, our primary interest often centers on whether the type of data present within a dataset wields considerable influence. We modify the definition of dataset’s influence so as to better align with this research question. Thus, for every training dataset $\mathcal{Z}$, we train a model for the same number of training steps $T$ instead of the entire dataset: $\bm{\theta}_{-\mathcal{Z}}=\mathrm{arg}\min_{\bm{\theta}}\sum_{t=1}^{T}L(z_{t},\bm{\theta})$ where $z_{t}\sim\mathcal{D}\setminus\mathcal{Z}$. This setup is practical for evaluating the influence of datasets of different sizes. [^1] In Appendix A.1, we show the conventional leave-dataset-out setup overestimates the influence of large datasets.
@@ 3.1 UnTrac
where $T$ is the number of unlearning steps. Here, $\bm{\theta}_{i}$ is updated via gradient ascent, which maximizes the loss of the $i$-th batch $z_{i}$ in the training dataset $\mathcal{Z}$. If we use stochastic gradient ascent, the updated parameters can be written as Equation (3); however, any optimizer can be used for unlearning.
@@ 3.2 UnTrac-Inv
This alternative influence can be regarded as an approximation to UnTrac, Equation (2). Note that $\bm{\theta}_{j}=\bm{\theta}_{j-1}+\eta_{j}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{j-1})$, and we use the first-order approximation $L(z_{i},\bm{\theta}_{j})-L(z_{i},\bm{\theta}_{j-1})\approx\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{j-1})(\bm{\theta}_{j}-\bm{\theta}_{j-1})$. Equation (2) and (4) can then be re-approximated as Equation (5) and (6), respectively.
@@ 4.1 TracIn, GradDot & GradCos
Pruthi et al. (2020) proposed TracIn, which traces the influence of a training example by the total change in test loss during training. While the original TracIn assesses the influence of individual training examples, it can be easily extended to assess a whole dataset. As accounting for the loss reduction at every step is computationally expensive, they approximate it using model checkpoints, assuming that every example in the training dataset is encountered once between these checkpoints. By approximating the loss reduction with gradients at each checkpoint $t$: $L(z^{\prime}_{j},\bm{\theta}_{t})-L(z^{\prime}_{j},\bm{\theta}_{t+1})\approx\eta_{t}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{t})^{\top}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{t})$, TracIn is defined as:
@@
where $\mathcal{T}_{cp}$ denotes the set of training steps where the model checkpoints are saved. As using multiple checkpoints induces substantial overhead, only the last checkpoint is often used in practice Schioppa et al. (2023), which is referred to as GradDot. Barshan et al. (2020) pointed out that some outlier training examples have significantly large gradients, leading to an overestimation of their influences. Therefore, normalizing the gradients (i.e., replacing the dot product with cosine similarity) can be effective, referred to as GradCos Han and Tsvetkov (2021); Akyurek et al. (2022).
@@ 4.2 Hessian-based Influence Functions
Hessian-based influence functions (HIF) are grounded in robust statistics Hampel (1974); Cook and Weisberg (1982) and were introduced to deep learning by Koh and Liang (2017). Koh et al. (2019) used HIF to assess the influence of multiple training examples, and the estimated influence correlates well with the ground-truth. Given a trained model $\bm{\theta}_{0}$, HIF estimates the influence of a training dataset $\mathcal{Z}$ on a test dataset $\mathcal{Z}^{\prime}$ as Equation (LABEL:eq:influence_functions).
@@ 4.3 Connection to UnTrac & UnTrac-Inv
GradDot, GradCos, and HIF can be viewed as special cases of UnTrac and UnTrac-Inv. Suppose a single training batch contains all training examples. The increase in test loss after a single step of unlearning can be approximated as Equation (9).
$\approx\sum_{j=1}^{N^{\prime}}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{0})^{\top}(\bm{\theta}_{1}-\bm{\theta}_{0})$
When we use SGD, Equation (9) corresponds to GradDot by substituting $\bm{\theta}_{1}-\bm{\theta}_{0}=\eta_{0}\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})$. Similarly, It corresponds to GradCos if we use RMSProp or Adam: $\bm{\theta}_{1}-\bm{\theta}_{0}=\eta_{0}\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})/\|\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})\|$, and corresponds to HIF if we use Newton’s method: $\bm{\theta}_{1}-\bm{\theta}_{0}=\sum_{i=1}^{N}\bm{H}_{\bm{\theta}}^{-1}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})$. In the same way, these influence functions can also be regarded as a special case of UnTrac-Inv. HIF can be regarded as providing an approximation to UnTrac.
@@ 5 Experiments
We evaluate if our methods can measure the influence of training datasets across different model architectures (encoder-decoder and decoder-only) and training setups (pretraining and fine-tuning). [^2] The code and data are available at: https://github.com/misonuma/untrac/. More implementation details are noted in Appendix A.2.
Since we would not generally have a validation set (i.e., ground-truth attributions), it is impractical to tune the hyperparameters for each experiment. Thus, we set the hyperparameters based on the experimental results on Toxigen, one of the test datasets used in Section 5.2. The same hyperparameters are used across all the experiments. We use Adam Kingma and Ba (2014) with a constant learning rate of 5e-5, $\beta_{1}$ = 0.9, and $\beta_{2}$ = 0.999, while Adafactor is used in Section 5.1 due to memory constraints. Gradient clipping is turned off during unlearning. As for UnTrac, we set the batch size to 1 and run unlearning for 1 epoch. As discussed in Section 3.2, UnTrac-Inv requires a small number of unlearning steps and a large batch size. Hence, we set the batch size as 256 (a single batch contains all test examples) and the number of unlearning steps (epochs) as 5. In Section 6, we discuss the hyperparameter sensitivity of our methods.
@@
Following Koh and Liang (2017), we approximate the inverse Hessian by LISSA Agarwal et al. (2017), where the number of iterations is set to 10. We also use Arnoldi iteration with low-rank eigenvector projection following Schioppa et al. (2022). We set the number of Arnoldi iterations to $n$ = 25 and the number of eigenvectors to $\tilde{p}$ = 25 in Section 5.1 due to the memory constraints, while $n$ = 200 and $\tilde{p}$ = 100 in Section 5.2. Following the previous studies, the gradients of training datasets are normalized, and 256 training examples are used to approximate the Hessian of training loss.
Synthetic Datasets A Input Output
Test (Task P/Format P) What is the number that comes after {0}? {1}
Train1 (Task P/Format P) Determine the number that succeeds {two}. Provide your answer in numerical form. {3}
Train2 (Task P/Format Q) Determine the number that succeeds {one}. Provide your answer in words. {two}
Train3 (Task Q/Format P) Determine the length of ‘{problem}’. Provide your answer in numerical form. {7}
Train4 (Task Q/Format Q) Determine the length of ‘{align}’. Provide your answer in words. {five}
Synthetic Datasets B Input Output
Test (Task P/Format P) What letter remains in ‘{xmais}’ after extracting ‘{s}’, ‘{x}’, ‘{m}’, ‘{i}’? {a}
Train1 (Task P/Format P) Identify the character left after removing ‘{0}’, ‘{1}’, ‘{4}’, ‘{7}’ from ‘{71b40}’. {b}
Train2 (Task P/Format Q) Identify the character left after removing ‘{6}’, ‘{7}’, ‘{5}’, ‘{2}’ from ‘{27516}’. {1}
Train3 (Task Q/Format P) Identify the part of speech of ‘{problem}’. Select your answer with the associated letter. Choices: a. noun b. verb c. adjective d. adverb {a}
Train4 (Task Q/Format Q) Identify the part of speech of ‘{align}’. Select your answer with the associated number. Choices: 0. noun 1. verb 2. adjective 3. adverb {1}
Table 1: Example of the synthetic datasets A (top) and B (bottom). {Strings in braces} are varied with each example.
Synthetic Datasets A Synthetic Datasets B
Train Dataset 1(TaskP/FmtP) 2(TaskP/FmtQ) 3(TaskQ/FmtP) 4(TaskQ/FmtQ) 1(TaskP/FmtP) 2(TaskP/FmtQ) 3(TaskQ/FmtP) 4(TaskQ/FmtQ)
GradDot 0.497 0.930 -1.680 0.253 -0.561 1.730 -0.655 -0.513
GradCos 0.993 0.573 -1.637 0.071 1.197 0.708 -1.307 -0.598
HIF (Arnoldi) 1.059 0.566 -0.026 -1.599 1.368 0.015 0.074 -1.457
HIF (LISSA) 0.156 -1.674 0.658 0.860 -0.600 1.708 -0.785 -0.323
TracIn 0.090 1.575 -1.083 -0.582 -0.174 1.638 -1.060 -0.404
UnTrac 1.443 0.344 -0.574 -1.212 1.492 0.175 -0.414 -1.253
UnTrac-Inv 0.980 1.014 -1.104 -0.891 1.056 0.905 -0.716 -1.245
Ground Truth 1.263 0.692 -1.064 -0.891 1.150 0.837 -1.025 -0.962
Table 2: Influence of the training datasets estimated by each method. The average is shown across four runs. The values are standardized for each method to normalize the range of values. “Fmt” is the abbreviation of “Format”.
@@ 5.1 Tracing Influential Training Tasks
Dataset
We create two synthetic datasets, each containing one test dataset and four training datasets. The test datasets contain examples of task P with the output format P (Task P/Format P). The task and output format of training datasets are set with respect to the test dataset as follows:
  1. 1.
    Training dataset 1 (Task P/Format P): similar task P with the same output format P.
  2. 2.
    Training dataset 2 (Task P/Format Q): similar task P with different output format Q.
  3. 3.
    Training dataset 3 (Task Q/Format P): irrelevant task Q with the same output format P.
  4. 4.
    Training dataset 4 (Task Q/Format Q): irrelevant task Q with different output format Q.
Table 1 presents examples of the synthetic datasets A and B. Each training dataset and test dataset consists of 256 examples. The model is trained for 512 steps on the mixture of the four training datasets with a batch size of 2.